Solve the system by addition
5x-4y=-5
x-2y=-7
Please help solve this

Answer by checkley75(3666)

5X-4Y=-5 NOW MULTIPLY THE SECOND EQUATION BY -5 AND ADD THEM:
-5X+10Y=35
---------------------------------
6Y=30
Y=30/6
Y=5 ANSWER. NOW SUBSTITUTE 5 FOR Y AND SOLVE FOR X THUS:
5X-4*5=-5
5X-20=-5
5X=-5+20
5X=15
X=15/5
X=3 ANSWER
PROOF
3-2*5=-7
3-10=-7
-7=-7